Chapter 17: Financial Statement Analysis OBJ,3,4,5 PR 17-4A Measures of liquidity, solvency, and profitability h receivables, 18.3 Marshall common stock was $82.60 on December 31, 20Y2 5. Number of days sales in receivablahe cmparative financial statements of Manshall Inc. are as follows. The market price of Marshall Ine Comparative Retained Earnings Statement For the Years Ended December 31, 20Y2 and 20Y1 $3704,000 3264,000 600,000 $50.000 $4304000 $3.814000 $ 10,000 10,000 Retained earnings, January 1 Total Dividends: On preferred stock.. On common stock $ 110,000 110.000 54,194000 $3,704,000 Total dividends Retained earnings, December 31. Marshall Inc. Comparative Income Seatement Years Ended December 31, 20Y2 and 201 For the 10.850,000 $10,000.000 6.000,000 5450.000 Selling expenses Administrative expenses.. Total operating expenses income from operations 2,170000 $ 2000,000 16275001500,000 3.797.500 5 3,500,000 $ 1,052,500 $1,050,000 99.500 20,000 $ 1,152.000 % 1,070,000 132000 120.000 1020000 950,.000 Other expense (interest) Income before income tax Income tax expens Net income S 600000 $50.000 Marshall Inc. Comparative Balance Sheet December 31, 20Y2 and 20Y1 Current assets $1,050,000 950,000 01,000 585,000 Marketable securities 420,000 420,000 500000 380,000 Inventories Prepaid expenses Total current assets. 800,000 800,000 $9,024.000 $8.254.000 S 880000 800,000 Total assets Long-term Sabilities: Mortgage note payable 6% . . . Bonds payable, 4% Total liabilities Preferred 4% snock.SS par Common stock, Spar 3,000.000 $3,200.0003000,000 $4080,000$3.800,000 Total long-term Eabilities 250000 $ 250.000 500,000 3.704,000 944,000 $4,454,000 9.024,000$8254,000 .. Retained earnings Total stockholders equity 500,000 4,194,000 Continued)
media%2F589%2F589b0163-aff8-4799-8228-61
Answer #1

1. Working Capital = Currents Assets - Current Liability

= $2,464,000 - $880,000

= $1,584,000

( current assets & current liabilities amount is given in comparative balance sheet in question)

2. Current Ratio = Current Assets / Current Liabilities

= 2,464,000 / 880,000

= 2.8

3. Quick Ratio = (Total Current Assets - Inventory - Prepaid expenses) / Current Liabilities

= (2,464,000 - 420,000 - 108,000) / 880,000

= 2.2

4. Account receivable turnover = Net Credit Sales / Average account receivable

Net Credit sales (given in income statement in question above (sales) ) = $10,850,000

Average account receivable = opening account receivable + closing account receivable / 2

= 500,000 + 585,000 / 2

= $542,500

therefore,

Account receivable turnover = 10,850,000 / 542,500

= 20

6. Inventory Turnover = Cost of Goods Sold / Average Inventory

Cost of Goods Sold = $6,000,000 (given in income staatement)

Average Inventory = Opening Inventory + Closing Inventory / 2 (opening and closing given in balance sheet)

= 380,000 + 420,000 / 2

= 400,000

Therefore,

Inventory turnover = 6,000,000 / 400,000

= 15

8. Ratio of Fixed Assets to long term liability = Fixed Asset / Long term liability

Fixed Assets = Total Assets - Total current assets (Amounts given in the balance sheet)

= 9,024,000 - 2,464,000

= 6,560,000

Long Term liability = 3,200,000 ( given in balance sheet)

Therefore,

Ratio of Fixed Assets to long term liability = 6560000 / 3200000

= 2.05

= 2.1 (rounded off to one decimal place)

9. Ratio of liabilities to stock holders equity = Total liabilities / total stockholder's equity

(this is also called debt to equity ratio)

= 4,080,000 / 4,944,000

= 0.825

=0.8 (rounded off to one decimal place)

> The answer is incomplete

Simon Young Thu, Nov 11, 2021 4:42 PM